OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (harsimran_hs4):

since the glacier is melting at a constant rate so , total time = 2002 - 1924 = 78 years total decrease = 6 - 1 = 5 acres rate of decrease = total decrease / total time = 5/ 78 acres / year

OpenStudy (anonymous):

im lose

OpenStudy (harsimran_hs4):

go step by step and tell where the problem is

OpenStudy (anonymous):

total time = 2002 - 1924 = 78 years total decrease = 6 - 1 = 5 acres rate of decrease = total decrease / total time = 5/ 78 acres / year

OpenStudy (anonymous):

got how im lost

OpenStudy (harsimran_hs4):

total time = 2002 - 1924 = 78 years clear ?

OpenStudy (anonymous):

the linear equation should be something close to y=-8/125x + 6 :)

OpenStudy (anonymous):

is that right harsmran?

OpenStudy (harsimran_hs4):

haven`t solved it before hand.. i do it alongside so no idea right now

OpenStudy (anonymous):

I just did a rough calculation, but it should be something similar to what I have written.

OpenStudy (harsimran_hs4):

@sexybadass22 did you get the first step about years ?

OpenStudy (anonymous):

78

OpenStudy (anonymous):

A certain mountain has a elevation of 19,465 feet in 1924 , the glasier on this peak covered 6 acres . by 2002 this glasier had melted to only 1 acre . (a)assume that this glasier melted at a contant rate each year (b)use your answer from part a to write a liner equation that gives the acreage a of this glasier t years past 1924

OpenStudy (anonymous):

the height of the mountain is of no consequence for this question.